# Measuring What Makes You Unique: Difference-Aware User Modeling for Enhancing LLM Personalization

### v1

@@ Appendix B Baselines Details
Non-Perso:: This method does not utilize any user-specific information to assist in personalized review generation. The model’s input consists only of the item’s title and description, along with the rating and title of the review to be generated.
@@ Appendix C Evaluation Metrics
| Datasets () | Books | Movies & TV | CDs & Vinyl | |||
| Methods () | R-L | BLEU | R-L | BLEU | R-L | BLEU |
| Non-Perso | 0.1438 | 1.2313 | 0.1262 | 0.3510 | 0.1262 | 0.4880 |
| RAG | 0.1617 | 4.1365 | 0.1414 | 1.7926 | 0.1413 | 2.0186 |
| IntSum | 0.1606 | 4.6217 | 0.1410 | 2.0803 | 0.1414 | 2.4589 |
| LLM-TRSR | 0.1616 | 5.0839 | 0.1408 | 2.1157 | 0.1416 | 2.4600 |
| CICL | 0.1654 | 5.3592 | 0.1422 | 2.1648 | 0.1412 | 2.3713 |
| Persona-DB | 0.1670 | 5.4166 | 0.1411 | 2.0175 | 0.1408 | 2.2963 |
| DPL | 0.1708 | 5.9593 | 0.1454 | 2.4065 | 0.1446 | 2.7514 |
| Methods () | R-1 | R-L | METEOR | BLEU | S-72B | S-GPT |
| Non-Perso | 0.3016 | 0.1508 | 0.1930 | 2.5779 | 0.6132 | 0.4650 |
| RAG | 0.3342 | 0.1765 | 0.2777 | 6.5834 | 0.6233 | 0.4833 |
| IntSum | 0.3252 | 0.1744 | 0.2851 | 6.4804 | 0.6278 | 0.4801 |
| LLM-TRSR | 0.3210 | 0.1718 | 0.2813 | 6.0386 | 0.6334 | 0.4890 |
| CICL | 0.3232 | 0.1679 | 0.2727 | 5.5111 | 0.6259 | 0.4890 |
| Persona-DB | 0.3252 | 0.1689 | 0.2768 | 5.6887 | 0.6183 | 0.4814 |
| DPL | 0.3366 | 0.1818 | 0.2879 | 6.6967 | 0.6461 | 0.5142 |
| Datasets () | Books | Movies & TV | CDs & Vinyl | ||||||||||
| Methods () | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | |
| Non-Perso | 0.2683 | 0.1512 | 0.5079 | 0.4360 | 0.2275 | 0.1304 | 0.4767 | 0.3242 | 0.2392 | 0.1336 | 0.4701 | 0.3409 | |
| RAG | N=1 | 0.3067 | 0.1820 | 0.5211 | 0.4114 | 0.2535 | 0.1533 | 0.4208 | 0.2476 | 0.2649 | 0.1553 | 0.4414 | 0.2989 |
| N=2 | 0.3167 | 0.1997 | 0.5609 | 0.4429 | 0.2682 | 0.1665 | 0.4712 | 0.2975 | 0.2755 | 0.1654 | 0.4803 | 0.3320 | |
| N=4 | 0.3220 | 0.2084 | 0.5987 | 0.4650 | 0.2785 | 0.1778 | 0.5117 | 0.3307 | 0.2880 | 0.1785 | 0.5158 | 0.3560 | |
| IntSum | N=1 | 0.3089 | 0.1950 | 0.5174 | 0.3628 | 0.2646 | 0.1677 | 0.3991 | 0.2079 | 0.2725 | 0.1653 | 0.3985 | 0.2391 |
| N=2 | 0.3172 | 0.2131 | 0.5703 | 0.4473 | 0.2758 | 0.1801 | 0.4798 | 0.2820 | 0.2854 | 0.1780 | 0.4888 | 0.3301 | |
| N=4 | 0.3188 | 0.2238 | 0.6076 | 0.4575 | 0.2821 | 0.1904 | 0.5368 | 0.3308 | 0.2932 | 0.1909 | 0.5364 | 0.3608 | |
| LLM- TRSR | N=1 | 0.3107 | 0.2024 | 0.5817 | 0.4442 | 0.2682 | 0.1747 | 0.5047 | 0.3018 | 0.2797 | 0.1747 | 0.5007 | 0.3431 |
| N=2 | 0.3146 | 0.2170 | 0.5975 | 0.4505 | 0.2749 | 0.1825 | 0.5289 | 0.3186 | 0.2883 | 0.1845 | 0.5271 | 0.3537 | |
| N=4 | 0.3174 | 0.2228 | 0.6164 | 0.4606 | 0.2803 | 0.1906 | 0.5557 | 0.3413 | 0.2931 | 0.1927 | 0.5413 | 0.3667 | |
| CICL | N=1 | 0.3157 | 0.2013 | 0.4921 | 0.3464 | 0.2652 | 0.1356 | 0.1689 | 0.2044 | 0.2741 | 0.1666 | 0.3921 | 0.2319 |
| N=2 | 0.3243 | 0.2162 | 0.5659 | 0.4221 | 0.2761 | 0.1798 | 0.4767 | 0.2804 | 0.2879 | 0.1805 | 0.4777 | 0.3092 | |
| N=4 | 0.3230 | 0.2282 | 0.5918 | 0.4669 | 0.2834 | 0.1889 | 0.5236 | 0.3271 | 0.2944 | 0.1805 | 0.5197 | 0.3488 | |
| Persona -DB | N=1 | 0.3184 | 0.2047 | 0.5281 | 0.4069 | 0.2667 | 0.1679 | 0.4243 | 0.2381 | 0.2731 | 0.1653 | 0.4115 | 0.2538 |
| N=2 | 0.3249 | 0.2228 | 0.5987 | 0.4631 | 0.2748 | 0.1773 | 0.4904 | 0.2984 | 0.2842 | 0.1773 | 0.4887 | 0.3239 | |
| N=4 | 0.3264 | 0.2305 | 0.6202 | 0.4902 | 0.2812 | 0.1882 | 0.5377 | 0.3321 | 0.2911 | 0.1867 | 0.5357 | 0.3623 | |
| DPL | N=1 | 0.3175 | 0.2010 | 0.5338 | 0.4069 | 0.2726 | 0.1731 | 0.4408 | 0.2542 | 0.2855 | 0.1761 | 0.4522 | 0.2829 |
| N=2 | 0.3256 | 0.2180 | 0.5823 | 0.4435 | 0.2858 | 0.1880 | 0.5117 | 0.3239 | 0.2995 | 0.1909 | 0.5162 | 0.3449 | |
| N=4 | 0.3270 | 0.2347 | 0.6095 | 0.4877 | 0.2896 | 0.1952 | 0.5493 | 0.3515 | 0.3059 | 0.2006 | 0.5532 | 0.3762 | |
@@
S-72B: S-72B is a metric that assesses the quality of generated text by comparing it to a reference text using the Qwen-2.5-72B-Instruct-AWQ model. We design the prompt following previous works Li et al. (2024c); Liu et al. (2023); Salemi et al. (2025) and make some reasonable modifications to better suit our task. In this evaluation, the prompt to the model includes the generated review, the user’s real review as the reference, as well as relevant metadata such as the review title, rating, and item details, including its title and description. We also perform scoring inference using vLLM, with the temperature set to 0.8 and top_p set to 0.95. The prompt is depicted in Figure C.
@@
Although traditional automated evaluation metrics such as ROUGE-L and BLEU Papineni et al. (2002) have been questioned for their ability to distinguish personalization from other subtle semantic aspects Koh et al. (2022), these metrics are still used in some studies Li et al. (2024a); Zhang et al. (2025a) and hold exploratory value. Therefore, we present a comparison of our method and the baseline methods based on ROUGE-L and BLEU, as shown in Table 5. The descriptions of these two metrics are provided below.
BLEU: BLEU (Bilingual Evaluation Understudy) is a metric that evaluates text similarity by measuring the precision of n-gram matches between the generated text and the reference text. In our experiments, we use SacreBLEU Post (2018), which is a standardized version of BLEU. The higher the BLEU score, the more similar the generated text is to the reference.
@@ D.1 Running Environments
We utilize the vLLM222https://github.com/vllm-project/vllm library Kwon et al. (2023) as the inference engine for LLMs in generating and the Sentence-Transformers333https://www.sbert.net/ Reimers and Gurevych (2019) library for LLMs in embedding in a Python 3.11.11 environment. All experiments are conducted on a single NVIDIA H100 NVL GPU with 96GB of GPU memory.
@@ D.2 Hyperparameter Configurations
We configure the model with a maximum output length of 2048 tokens while keeping the input length unrestricted. Following the settings in LongLaMP, we set the temperature to 0.8 and the parameter top_p to 0.95. These settings remain consistent across all experiments.
Appendix E Additional Experiments
E.1 Model of Different Size
@@
To investigate the generalization capability of our approach, we conduct experiments to evaluate its performance across models of different sizes. Specifically, we consider the Qwen2.5-32B-Instruct model. For each method, we utilize 8 retrieved historical data points. To reduce resource costs, we perform our experiments on the Books category only and report the results in ROUGE-1, ROUGE-L, METEOR, BLEU, S-72B, and S-GPT metrics. The experimental results are shown in Table 6, from which the following observations can be drawn:
@@
When using the 32B model, its more powerful capabilities lead to performance improvements across almost all methods compared to the 14B model. However, DPL still achieves state-of-the-art performance, outperforming the best baseline metrics by an average of 2.26%. This demonstrates its strong generalization capability and effectiveness in the personalized review generation task.
@@
E.2 Number of Retrieved Reviews
We conducted experiments to explore the impact of different numbers of retrieved reviews on performance. Specifically, we evaluated cases with 1, 2, and 4 retrieved reviews. The experimental results presented in Table 7 lead to the following observations:
@@ E.2 Number of Retrieved Reviews
When the number of retrieved reviews is 4, our DPL method outperforms other baseline approaches on the CDs & Vinyl dataset and achieves competitive performance on the remaining two datasets. This demonstrates the superiority of our DPL method.
Appendix F Overview of Templates & Prompts
@@
Appendix G Case Studies
@@
G.1 Case Study: Comparison between Difference-aware Representation and User Profile Summary
@@
G.2 Case Study: Outputs of DPL Components

### v2

@@ Appendix B Baselines Details
Non-Perso: This method does not utilize any user-specific information to assist in personalized review generation. The model’s input consists only of the item’s title and description, along with the rating and title of the review to be generated.
@@ Appendix C Evaluation Metrics
BLEU: BLEU (Bilingual Evaluation Understudy) is a metric that evaluates text similarity by measuring the precision of n-gram matches between the generated text and the reference text. In our experiments, we use the SacreBLEU333https://github.com/mjpost/sacrebleu Post (2018) library, which is a standardized version of BLEU. The higher the BLEU score, the more similar the generated text is to the reference.
S-72B: S-72B is a metric that assesses the quality of generated text by comparing it to a reference text using the Qwen-2.5-72B-Instruct-AWQ model. We design the prompt following previous works Liu et al. (2023); Li et al. (2024b); Salemi et al. (2025) and make some reasonable modifications to better suit our task. In this evaluation, the prompt to the model includes the generated review, the user’s real review as the reference, as well as relevant metadata such as the review title, rating, and item details, including its title and description. To ensure robustness, we run the scoring process with two different random seeds and report the average of their results. We also perform scoring inference using the vLLM library, with the temperature set to 0.8 and top_p set to 0.95. The prompt is depicted in Figure C.
@@ D.1 Running Environments
We utilize the vLLM444https://github.com/vllm-project/vllm library Kwon et al. (2023) as the inference engine for LLMs in generating and the Sentence-Transformers555https://www.sbert.net/ Reimers and Gurevych (2019) library for LLMs in embedding in a Python 3.11.11 environment. All experiments are conducted on a single NVIDIA H100 NVL GPU with 96GB of GPU memory.
@@ D.2 Hyperparameter Configurations
We configure the model with a maximum output length of 2048 tokens while keeping the input length unrestricted. Following the settings in LongLaMP, we set the parameter top_p to 0.95. These settings remain consistent across all experiments.
Appendix E Efficiency Analysis
In this section, we briefly discuss the efficiency of our proposed DPL.
We evaluate the cost of selecting comparison users using the K-means method with pre-computed review embeddings. Since the number of involved users is small, the selection process takes only 0.002s per key historical review on average. In our main experiments, we use 8 key historical reviews, resulting in an additional cost of approximately 0.016s per generation, which is generally acceptable for each time generation.
Additionally, since our method uses a fixed number of comparison users for difference extraction, the computational cost of most components remains stable as the data size increases. The only exception is the comparison user selection process, which uses K-means clustering (). The cost of this step depends on the number of users who have interacted with a given item (denoted as ), with a complexity of , where t is the number of iterations. Importantly, this step is lightweight—clustering 10,000 samples takes under 0.3 seconds in our setup — compared to LLM inference. Thus, the additional overhead is minimal, and our method scales comparably to baseline approaches.
Appendix F Additional Experiments
F.1 Yelp Dataset
To further validate the generalizability and robustness of our proposed method, we conduct additional experiments on the Yelp666https://business.yelp.com/data/resources/open-dataset dataset. This dataset presents a different domain from our primary evaluation benchmarks, allowing us to assess whether the observed performance gains extend to a broader range of real-world scenarios.
Dataset Construction. The Yelp Open Dataset is a large-scale, publicly available collection of user-generated reviews, primarily covering local businesses such as restaurants, shopping centers, and service providers across various cities. It contains rich information, including user reviews, ratings, business metadata, and timestamps, making it a valuable resource for studying the personalized review generation task. Following the same protocol as applied to the Amazon dataset, we perform data filtering and cleaning to ensure consistency. In particular, we retain several key fields: user_id, which identifies the user who wrote the review; business_id, which specifies the reviewed business; business_name, the name of the business; stars, the rating score of the review; text, which contains the full content of the review; and date, which we represent by timestamp. The statistics of the processed Yelp dataset is shown in Table 5. Our processed Yelp dataset is also available on Huggingface777https://huggingface.co/datasets/SnowCharmQ/DPL-Yelp.
Implementation Details. We largely follow the same implementation setup as in the main experiments, except that the item description is removed from the prompt. in Equation (3) is set to 8, and in Equation (4) is set to 4. Due to the large scale of the original dataset, we randomly sample 100 data points from the full set in each run to save computational resources and time while maintaining reliability. This sampling process is repeated five times, and we report the mean and standard deviation of the results across the five runs.
Splits () | #data | Profile Size | Output Length |
Train | 229944 | 38.5641.75 | 681.81394.80 |
Val | 28743 | 43.0641.69 | 672.05394.69 |
Test | 28743 | 44.0641.69 | 659.71390.20 |
Methods () | R-1 | R-L | METEOR | BLEU |
Non-Perso | ||||
RAG | ||||
IntSum | ||||
LLM-TRSR | ||||
CICL | ||||
Persona-DB | ||||
DPL |
Experimental Results. We report the results on conventional metrics here and also perform significance testing to assess the statistical reliability of the observed differences. The results are shown in Table 6, from which we observe that our DPL method consistently achieves the best performance. Moreover, it demonstrates statistically significant improvements over all baselines across the four evaluation metrics.
Methods () | R-1 | METEOR | S-72B |
Non-Perso | |||
RAG | |||
IntSum | |||
LLM-TRSR | |||
CICL | |||
Persona-DB | |||
DPL |
F.2 Model of Different Size
@@
To further investigate the generalization capability of our approach, we conduct experiments to evaluate its performance across models of different sizes. Specifically, we consider the Qwen2.5-32B-Instruct model. For each method, we utilize 8 retrieved historical data points. To reduce resource costs, we perform our experiments on the Books category only and report the results in ROUGE-1, METEOR, and S-72B metrics. The experimental results are shown in Table 7, from which the following observations can be drawn:
@@
When using the 32B model, its more powerful capabilities lead to performance improvements across almost all methods compared to the 14B model. However, DPL still achieves state-of-the-art performance. This demonstrates its strong generalization capability and effectiveness in the personalized review generation task.
@@
Datasets () | Movies & TV | CDs & Vinyl | Books | |||||||
Methods () | R-1 | MET. | S-72B | R-1 | MET. | S-72B | R-1 | MET. | S-72B | |
Non-Perso | 0.2275 | 0.1304 | 0.5406 | 0.2392 | 0.1336 | 0.5379 | 0.2683 | 0.1512 | 0.5539 | |
RAG | N=1 | 0.2540 | 0.1540 | 0.4858 | 0.2636 | 0.1546 | 0.5082 | 0.3007 | 0.1784 | 0.5609 |
N=2 | 0.2670 | 0.1658 | 0.5315 | 0.2768 | 0.1660 | 0.5434 | 0.3158 | 0.1971 | 0.5836 | |
N=4 | 0.2802 | 0.1794 | 0.5617 | 0.2884 | 0.1787 | 0.5705 | 0.3211 | 0.2116 | 0.6145 | |
N=8 | 0.2845 | 0.1897 | 0.5929 | 0.2926 | 0.1876 | 0.5933 | 0.3206 | 0.2173 | 0.6196 | |
IntSum | N=1 | 0.2643 | 0.1682 | 0.4618 | 0.2735 | 0.1671 | 0.4582 | 0.3155 | 0.2030 | 0.5533 |
N=2 | 0.2757 | 0.1807 | 0.5305 | 0.2865 | 0.1809 | 0.5466 | 0.3179 | 0.2144 | 0.5975 | |
N=4 | 0.2835 | 0.1916 | 0.5752 | 0.2938 | 0.1921 | 0.5848 | 0.3167 | 0.2256 | 0.6151 | |
N=8 | 0.2845 | 0.1972 | 0.6073 | 0.2949 | 0.1968 | 0.6014 | 0.3201 | 0.2340 | 0.6265 | |
LLM- TRSR | N=1 | 0.2664 | 0.1728 | 0.5578 | 0.2811 | 0.1756 | 0.5602 | 0.3090 | 0.2015 | 0.5943 |
N=2 | 0.2760 | 0.1838 | 0.5816 | 0.2876 | 0.1835 | 0.5770 | 0.3125 | 0.2137 | 0.6057 | |
N=4 | 0.2808 | 0.1915 | 0.5942 | 0.2942 | 0.1941 | 0.5981 | 0.3169 | 0.2269 | 0.6177 | |
N=8 | 0.2822 | 0.1955 | 0.6178 | 0.2937 | 0.1969 | 0.6083 | 0.3143 | 0.2294 | 0.6334 | |
CICL | N=1 | 0.2658 | 0.1682 | 0.4538 | 0.2737 | 0.1656 | 0.4575 | 0.3124 | 0.1977 | 0.5331 |
N=2 | 0.2759 | 0.1787 | 0.5242 | 0.2883 | 0.1797 | 0.5306 | 0.3230 | 0.2176 | 0.5748 | |
N=4 | 0.2836 | 0.1881 | 0.5731 | 0.2949 | 0.1890 | 0.5732 | 0.3211 | 0.2278 | 0.6196 | |
N=8 | 0.2873 | 0.1969 | 0.6142 | 0.3010 | 0.1987 | 0.5990 | 0.3255 | 0.2362 | 0.6391 | |
Persona -DB | N=1 | 0.2653 | 0.1658 | 0.4695 | 0.2719 | 0.1639 | 0.4637 | 0.3162 | 0.2000 | 0.5451 |
N=2 | 0.2754 | 0.1786 | 0.5407 | 0.2843 | 0.1761 | 0.5479 | 0.3212 | 0.2161 | 0.5950 | |
N=4 | 0.2819 | 0.1882 | 0.5822 | 0.2899 | 0.1861 | 0.5887 | 0.3267 | 0.2336 | 0.6309 | |
N=8 | 0.2840 | 0.1954 | 0.6170 | 0.2941 | 0.1937 | 0.6048 | 0.3268 | 0.2423 | 0.6435 | |
DPL | N=1 | 0.2753 | 0.1747 | 0.5015 | 0.2859 | 0.1765 | 0.5158 | 0.3164 | 0.1989 | 0.5659 |
N=2 | 0.2858 | 0.1877 | 0.5630 | 0.2997 | 0.1917 | 0.5704 | 0.3202 | 0.2175 | 0.5905 | |
N=4 | 0.2907 | 0.1949 | 0.5854 | 0.3071 | 0.2014 | 0.5958 | 0.3269 | 0.2311 | 0.6158 | |
N=8 | 0.2938 | 0.2009 | 0.6245 | 0.3066 | 0.2058 | 0.6236 | 0.3315 | 0.2438 | 0.6435 |
F.3 Number of Retrieved Reviews
We conducted experiments to explore the impact of different numbers of retrieved reviews on performance. Specifically, we evaluated cases with 1, 2, 4, and 8 retrieved reviews. To save resources and time, we only report the results with metrics ROUGE-1, METEOR, and S-72B here. The experimental results, reported from a single run with decoding temperature set to 0.8, are presented in Table 8, leading to the following observations:
@@ F.3 Number of Retrieved Reviews
Across different numbers of retrieved reviews, our DPL method consistently achieves performance that is comparable to or even surpasses the best-performing baselines. Notably, when the number of retrieved reviews reaches 8, DPL steadily outperforms all other methods across all three datasets. This demonstrates the robustness of our method and its ability to effectively leverage richer user history for enhanced personalization.
Appendix G Overview of Templates & Prompts
@@
Appendix H Case Studies
@@
H.1 Case Study: Comparison between Difference-aware Representation and User Profile Summary
@@
H.2 Case Study: Outputs of DPL Components